# LightTact: A Visual–Tactile Fingertip Sensor for Deformation-Independent Contact Sensing

### v1

@@ Abstract
Contact often occurs without macroscopic surface deformation, such as during interaction with liquids, semi-liquids, or ultra-soft materials. Most existing tactile sensors rely on deformation to infer contact, making such light-contact interactions difficult to perceive robustly. To address this, we present LightTact, a visual-tactile fingertip sensor that makes contact directly visible via a deformation-independent, optics-based principle. LightTact uses an ambient-blocking optical configuration that suppresses both external light and internal illumination at non-contact regions, while transmitting only the diffuse light generated at true contacts. As a result, LightTact produces high-contrast raw images in which non-contact pixels remain near-black (mean gray value <3<3) and contact pixels preserve the natural appearance of the contacting surface. Built on this, LightTact achieves accurate pixel-level contact segmentation that is robust to material properties, contact force, surface appearance, and environmental lighting. We further integrate LightTact on a robotic arm and demonstrate manipulation behaviors driven by extremely light contact, including water spreading, facial-cream dipping, and thin-film interaction. Finally, we show that LightTact’s spatially aligned visual-tactile images can be directly interpreted by existing vision-language models, enabling resistor value reasoning for robotic sorting.
@@ I Introduction
Contact occurs throughout physical interactions between robots and their environments, for example, between fingertips and objects during manipulation or between feet and terrain during locomotion. To perform such interactions safely, precisely, and robustly, robots must be able to detect and localize contact [dahiya2009tactile, luo2025tactile]. Vision-based tactile sensors (VBTSs) have gained prominence in recent years due to their high spatial resolution and low cost [li2025classification, li2024vision, xin2025vision, he2025survey]. However, most VBTSs sense contact indirectly by observing macroscopic deformation of a soft surface [yuan2017gelsight, yamaguchi2016combining, lin20239dtact, ward2018tactip, zhang2022deltact]. This deformation-dependent mechanism fundamentally requires measurable indentation to generate informative tactile signal. In contrast, many real-world interactions involve light contacts that induce little or no surface deformation. Examples include interactions with liquids or semi-liquids, and gentle contacts with extremely soft or thin materials. Such scenarios are common in everyday tasks, such as spreading water on a tabletop, dipping facial cream, or detecting a sheet of food film, but they remain challenging for deformation-based tactile sensors.
Several sensors attempt to bypass deformation by using frustrated total internal reflection (TIR) to highlight the contact regions of a transparent medium [han2005low, shimonomura2016robotic, zhang2023tirgel, sun2025soft]. However, ambient illumination and reflections from non-contact regions can also enter the camera, seriously corrupting the highlighted contact signal. As a result, TIR-based sensors typically function reliably only with monochromatic objects or under dark, tightly controlled environments–conditions incompatible with robotic tactile perception in everyday, unstructured settings. Thus, building a tactile sensor that can robustly detect contact without relying on surface deformation remains an open challenge.
@@
To address this challenge, we introduce LightTact, a compact fingertip visual–tactile sensor that achieves direct, optics-based contact detection independent of surface deformation (Fig. 1). The key innovation is our dedicated optical configuration that entirely suppresses both external light and internal illumination at non-contact regions, while allowing only the diffuse light generated at true contacts to reach the camera. This design fundamentally changes how contact is encoded in the raw sensor image, enabling three core properties:
  • High-contrast raw images. LightTact captures raw images with naturally high-contrast contact representation: non-contact pixels remain near-black (intensity <2<2), while contacting regions appear relatively bright.
  • Robust pixel-level contact segmentation. Supported by the inherent optical isolation, for the first time, LightTact achieves pixel-level contact segmentation that is robust to material properties, contact forces, surface appearance, and environmental lighting.
  • Aligned visual–tactile multimodality. The bright pixels simultaneously encode the localized visual appearance of the object and the contact region of the sensor surface, producing a naturally aligned multimodal signal suitable for robot learning and high-level reasoning.
Leveraging these properties, we integrate LightTact into a robotic system and demonstrate manipulation capabilities that remain difficult or inaccessible for existing tactile sensors. These include reliable interaction with liquids, semi-liquids, and extremely soft materials, as well as fine-grained sorting of tiny objects by directly prompting LightTact images to vision–language models. Together, these results indicate that LightTact enables a new class of manipulation behaviors by providing direct, optics-based sensing of true physical contact.
In summary, our contributions are as follows:
  • We propose an ambient-blocking optical layout that theoretically isolates all external and internal illumination from non-contact regions, enabling intuitive contact visualization without relying on deformation.
  • We introduce LightTact, a compact fingertip visual-tactile sensor that achieves robust contact segmentation with minimal computation cost.
  • We systematically evaluate LightTact across diverse materials, contact areas, and lighting conditions, and report its success and failure cases.
  • We demonstrate that LightTact unlocks manipulation capabilities that require reliable detection of light contact, and supports multimodal reasoning by directly prompting LightTact images to vision–language models.
@@ II-A Contact Sensing with Vision-Based Tactile Sensors
Vision-based tactile sensors (VBTSs) use internal cameras to capture visualized tactile signals, offering high spatial sensitivity at low cost. However, most existing VBTSs detect contact indirectly by observing macroscopic deformation of a soft sensing surface. Common approaches include tracking painted markers [yamaguchi2016combining, ward2018tactip, alspach2019soft, cui2021hand, zhang2022deltact, zhang2022tac3d], observing magnified reflective patterns [saga2007high, saga2013precise], or measuring luminance variations of reflective layers [yuan2017gelsight, lin20239dtact]. Fundamentally, these approaches require visible surface deformation, which imposes three constraints: the surface must be soft, the contacting object must be relatively harder, and the applied force must exceed a detectable deformation threshold. These requirements not only fix the sensor to a soft surface, but also severely limit its ability to perceive light contacts that do not generate measurable indentation, such as touching water, facial cream, or thin films.
To decouple sensing from mechanical deformation, several sensors [han2005low, shimonomura2016robotic, zhang2023tirgel, sun2025soft] apply total internal reflection (TIR) within a transparent medium. When an object contacts the medium, local frustration of TIR occurs, reflecting additional light toward the camera and thereby highlighting contact regions. However, this highlighting strategy is highly sensitive to ambient light. External light, reflections from non-contact portions of the object, and internal LED illumination can all enter the camera through non-contact regions, interfering with or overwhelming the intended TIR-based contact signal. As a result, these systems operate reliably only with monochromatic objects or under controlled, dark environments.
In contrast, LightTact completely suppresses both external and internal illumination at non-contact regions while capturing only the diffuse light generated at true contacts. This ensures that non-contact pixels remain near-black in raw images, enabling, for the first time, robust pixel-level contact segmentation across diverse lighting conditions and object appearances.
Similar objectives of reducing interference from external light have appeared in non-robotic optical devices, such as the Wyman–White fingerprint imager [wyman1965method] and the Drawing Prism [greene1985drawing]. However, these systems were designed for static fingerprint capture or artistic tracing, where requirements are substantially more relaxed. They do not enforce strict darkness of non-contact pixels, assume controlled lighting conditions, or restrict the types of interacting objects. As a result, they rely on bulky geometries, dark environments, and even a thin liquid layer to eliminate air gaps. Such conditions are incompatible with robotic tactile perception, where sensors must operate reliably in dynamic, unconstrained daily environments.
@@ III Methodology
Our goal is to design LightTact as a robotic tactile sensor capable of reliable, pixel-level contact segmentation across diverse objects and lighting conditions. To this end, we first introduce a novel optical layout and sensing principle and analyze its theoretical behavior in Section III-A. We then present how LightTact realizes this principle within a compliant, fingertip-scale form factor suitable for robotic integration in Section III-B. Following, we describe the key fabrication steps for constructing LightTact in Section III-C. Finally, we introduce our algorithms for contact segmentation in Section III-D.
@@ III-A Optical Layout and Sensing Principle
Figure 2: Optical layout and sensing principle of LightTact. (a) Core components: a transparent medium with a touching and viewing surface, an internal LED, and a camera. (b) External light entering through non-contact regions is rejected by refraction followed by total internal reflection (TIR). (c) Internal LED illumination reflecting from non-contact regions is also rejected by TIR when the LED is placed within the purple area to satisfy $\theta_{it}<\theta_{i}$. (d) At true contact, the air gap is removed and the contacting surface produces diffuse reflection. A subset of these rays reaches the viewing surface with incidence $\theta_{c}$ and refracts toward the camera.
Figure 2 (a) shows the core optical elements of LightTact: a transparent medium that interacts with external objects, an internal LED for illumination, and a camera that observes light exiting the medium. Two surfaces of the medium play distinct optical roles: the touching surface, where contact occurs, and the viewing surface, through which light exits toward the camera. The remaining surfaces are coated matte-black to absorb any light that reaches them. At the touching surface, combining two light sources (external light and internal illumination) with two region types (contact vs. non-contact) produces three relevant light behaviors for non-luminous objects: (1) external light incident on non-contact regions, (2) internal illumination incident on non-contact regions, and (3) internal illumination scattered at contact regions.
Our design objective is to obtain deformation-independent contact segmentation directly from the raw image. To achieve this, the optical layout must completely suppress both external and internal light at all non-contact regions, while transmitting only light generated at true contacts. A central mechanism enabling this property is the nonparallel wedge geometry: the touching and viewing surfaces form an angle $\theta_{tv}$ rather than being parallel. This geometry forces the three classes of light above to follow fundamentally different optical paths, allowing only the contact-scattered light to transmit through the viewing surface. Table I summarizes these behaviors, and the following paragraphs explain each in detail.
@@
Internal Contact Diffuse reflection Refraction $\theta_{tv}<\frac{\pi}{2}+\theta_{c}$
External light rejection at non-contact regions
As illustrated in Fig. 2 (b), external rays approach the touching surface from a wide range of directions. The sensing region corresponds to the span between the two depicted ray bundles. When external light enters the medium through non-contact areas, it refracts into the medium at angles below the critical angle $\theta_{c}$ due to the air–medium index difference. A portion of these refracted rays then propagate toward the viewing surface. By choosing a sufficiently large wedge angle ($\theta_{tv}>2\theta_{c}$), all such rays arrive at the viewing surface with incidence $\theta_{c}$, causing them to undergo total internal reflection (TIR) and redirecting them toward the black surfaces. Consequently, no external light from non-contact regions reaches the camera.
Internal illumination rejection at non-contact regions
Internal LED illumination should contribute to the image only at true contacts. As shown in Fig. 2(c), LED rays that reflect specularly off non-contact regions of the touching surface, must also reach the viewing surface with incidence $\theta_{c}$ to ensure TIR. This is guaranteed by constraining the LED position: the LED must lie entirely within the purple admissible region such that all emitted rays satisfy $\theta_{it}<\theta_{i}=\theta_{tv}-\theta_{c}$. Under this placement rule, every specularly reflected LED ray is rejected by TIR at the viewing surface, ensuring that no internal light leaks into the camera through non-contact regions.
@@
When an object touches the medium, the air gap disappears and the contacting surface produces diffuse reflection of the LED illumination (Fig. 2(d)). This generates a wide angular spread of rays, which can reach the viewing surface with incidence $\theta_{sv}\in\left(\theta_{tv}-\frac{\pi}{2},\frac{\pi}{2}\right)$. To observe the contacting surface, a subset of these scattered rays must refract through the viewing surface and reach the camera. This requires that the incidence range $\left(\theta_{tv}-\frac{\pi}{2},\frac{\pi}{2}\right)$ intersects the transmissible range $\left(-\theta_{c},\theta_{c}\right)$, which is satisfied when $\theta_{tv}<\frac{\pi}{2}+\theta_{c}$. Under this condition, a subset of diffuse-reflection rays transmits within the green regions and enters the camera, revealing the natural local appearance of the object. Meanwhile, all non-contact pixels remain black, producing a clean, high-contrast raw image that enables deformation-independent contact segmentation.
Our optical layout ensures that no light from non-contact regions reaches the camera by enforcing the condition $\theta_{tv}>2\theta_{c}$, which guarantees that all such rays are rejected by the viewing surface under TIR. Strictly speaking, $\theta_{tv}>2\theta_{c}$ is not the only possible configuration. If $\theta_{tv}<2\theta_{c}$, the camera could, in principle, be positioned so that any rays refracted through the viewing surface from non-contact regions fall outside its field of view. However, this alternative requires placing the camera significantly farther from the touching surface and introduces additional complications: rays refracted through the viewing surface may require additional black surfaces to prevent unintended reflections back toward the camera. We therefore adopt the simpler and more robust condition $\theta_{tv}>2\theta_{c}$ as our default configuration. A detailed analysis of camera-placement alternatives and their limitations is provided in Appendix V-A.
@@ III-B Sensor Design
Figure 3: Design of LightTact. (a) Assembled view of LightTact. (b) Components of LightTact illustrated in exploded view. (c) Schematic diagram of LightTact.
As shown in Fig. 3 (a), LightTact is extremely compact, with a fingertip-scale size of only $34.5\text{mm}\times 18\text{mm}\times 12\text{mm}$. Figures 3 (b) and (c) illustrate the internal assembly, which mirrors the optical system described in Section III-A: a transparent medium (2, 3), an internal LED (5), and a camera (7), all housed within a rigid shell (8), a surrounding black gel layer (1), and lightweight mounts (4, 6). The following paragraphs describe each component in detail.
Transparent gel and acrylic window. For most robotic applications, the touching surface should be soft to enable compliant physical interaction. Importantly, this softness is a design preference rather than a sensing requirement, since our optical principle does not rely on macroscopic deformation. However, a fully soft transparent medium would cause the viewing surface to deform under load, producing image distortion. To prevent this, the transparent medium combines a soft transparent silicone gel (for soft touching surface) with a rigid acrylic window (for rigid viewing surface). Both materials have similar refractive indices ($\approx 1.45$), which yields a critical angle of $\theta_{c}\approx 43^{\circ}$. To satisfy the conditions derived in Section III-A, namely $\theta_{tv}>2\theta_{c}\approx 86^{\circ}$ for suppressing all non-contact light, and $\theta_{tv}<90^{\circ}+\theta_{c}\approx 133^{\circ}$ for transmitting contact-scattered light, we set the touching and viewing surfaces to be perpendicular.
LED and camera. Internal illumination is provided by a single 2835-SMD LED. The LED is mounted at an oblique angle (approximately $45^{\circ}$) to provide more uniform lighting across the touching surface. Imaging is performed by a compact UVC camera (based on OV5693) with a $120^{\circ}$ field of view, placed close to the transparent medium to maintain the overall compactness of the sensor. Both the LED and the camera are fastened with lightweight mounts, which attach to the sensor shell using M2 screws and heat-set inserts.
Sensor shell. The sensor shell not only fixes the optical components in their required poses, but also constrains three critical geometric–optical properties: the transparent gel’s shape, the LED’s illumination range, the camera’s effective viewing range. These design aspects are summarized below.
@@
Although the sensor shell is built from black material to approximate the ideal “light-absorbing” boundaries assumed in Section III-A, real materials cannot completely absorb incident light. As indicated in Fig. 3 (c), the black surface labeled A may still reflect light toward the camera through non-contact regions. To suppress this effect, surface A is designed to be vertical ($\theta_{s}=\frac{\pi}{2}$), which minimizes the likelihood of specular reflections coupling into the viewing path. For certain robotic applications that require entering narrow spaces, the sensor may adopt a wedge-shaped profile. We analyze the configuration $\theta_{s}<\frac{\pi}{2}$ and its optical consequences in detail in Appendix V-C.
@@
To prevent the LED from directly illuminating surface A, the shell incorporates two baffles (surfaces B and C) that restrict the emission range of the LED. Although taller baffles improve shielding, their height is limited for two reasons: 1) If surface A extends above point D, the protruding inclined region will be directly illuminated and significantly increase the amount of light entering the camera via non-contact surface. 2) If surface B exceeds point E, it blocks a portion of the diffuse-reflection light from contact regions of the left side of the touching surface.
@@
The shell further restricts the camera’s field of view so that only the touching surface is observable. If the camera could see outside this region, additional stray light, such as the scattered light from the rough top or bottom surfaces of the acrylic window, could interfere the pixels corresponding to the touching surface.
Black gel. To preserve compliance over the entire contact area, the soft black gel is added around the perimeter of the transparent gel, forming a smooth transition between the touching surface and the rigid shell. It also covers the top surface of the acrylic window, which cannot function as part of the touching surface due to its rigidity and roughness.

### v2

@@ Abstract
Contact often occurs without macroscopic surface deformation, such as during interaction with liquids, semi-liquids, or ultra-soft materials. However, most existing tactile sensors rely on deformation to infer contact, making such light-contact interactions difficult to perceive robustly. To address this, we present LightTact, a visual-tactile fingertip sensor that makes contact directly visible via a deformation-independent principle. LightTact features an ambient-blocking optical configuration that suppresses both external light and internal illumination at non-contact regions, while transmitting only the scattered light generated at true contacts. As a result, LightTact produces high-contrast raw images in which non-contact pixels remain near-black (mean gray value <3<3) and contact pixels preserve the natural appearance of the contacting surface. Built on this, LightTact achieves accurate pixel-level contact segmentation that is robust to material properties, contact force, surface appearance, and environmental lighting. We further demonstrate that LightTact unlocks new robotic manipulation behaviors that require detection of extremely light contact, including water spreading, facial-cream dipping, and soft thin-film interaction. In addition, we show that LightTact’s spatially aligned visual-tactile images can be directly interpreted by vision-language models.
@@ I Introduction
Tactile sensing is essential for robotic manipulation, where robots must precisely regulate when, where, and how they physically interact with objects through contact [dahiya2009tactile, luo2025tactile]. In recent years, vision-based tactile sensors (VBTSs) have gained prominence due to their high spatial resolution and low cost [li2025classification, li2024vision, xin2025vision, he2025survey]. However, most VBTSs sense contact indirectly by observing macroscopic deformation of a soft surface [yuan2017gelsight, yamaguchi2016combining, lin20239dtact, ward2018tactip, zhang2022deltact]. This deformation-dependent mechanism fundamentally requires measurable indentation to generate an informative tactile signal. In contrast, many real-world interactions involve light contacts that induce little or no surface deformation. Examples include interactions with liquids or semi-liquids, and gentle contacts with extremely soft or thin materials. Such scenarios are common in everyday tasks, such as spreading water on a tabletop, dipping facial cream, or detecting a sheet of food film, but they remain challenging for deformation-based tactile sensors.
To relax this constraint, several sensors attempt to bypass deformation by using frustrated total internal reflection (TIR) to highlight the contact regions of a transparent medium [han2005low, shimonomura2016robotic, zhang2023tirgel, sun2025soft]. However, ambient illumination and reflections from non-contact regions can also enter the camera, seriously corrupting the highlighted contact signal. As a result, TIR-based sensors typically function reliably only with monochromatic objects or under dark, tightly controlled environments—conditions incompatible with robotic tactile perception in everyday, unstructured settings. Therefore, designing a tactile sensor that can robustly detect contact without relying on surface deformation remains an open challenge.
@@
In this work, we introduce LightTact, a camera-based fingertip sensor that makes physical contact directly visible in the raw image. LightTact enforces a bijective contact-visibility relationship: a pixel is visible if and only if it is in true physical contact. This enables robust pixel-level contact segmentation across liquids, ultra-soft materials, and rigid objects under diverse object appearances and bright ambient lighting (Fig. 1).
The key enabler is our ambient-blocking optical configuration that can, in theory, completely suppress both external light and internal illumination at non-contact regions, while transmitting only contact-generated scattered light to the camera through a transparent medium. We realize this behavior by adopting a side-view imaging optical layout with dedicated camera-LED-medium configurations, which jointly leverage refraction, specular reflection, and total internal reflection (TIR) to block non-contact light paths. To translate this principle into a practical robot fingertip, we co-design LightTact across optics, structure, materials, electronics, and fabrication. Finally, LightTact features the following properties:
  • High-contrast raw images where non-contact pixels remain near-black (gray value <3<3) while contact pixels preserve the surface appearance.
  • Pixel-level contact segmentation with lightweight computation, robust to material properties, contact force, ambient lighting, and surface appearance.
  • Aligned visual–tactile signals where both local appearance of the object and contact region of the sensor surface are naturally co-registered in the same RGB image.
  • Compact and compliant. LightTact has a fingertip-scale size of $12\times 18\times 34.5\penalty 10000\ \text{mm}$ and a soft surface, enabling direct integration into grippers and dexterous hands.
  • Affordable and reproducible. The primary cost is the RGB camera; the remaining components can be fabricated with standard tools for under $\mathdollar 20$.
  • Open-source. We will release the full hardware and software, along with a step-by-step fabrication tutorial.
Leveraging these properties, we integrate LightTact into a robotic system and demonstrate manipulation behaviors that remain difficult or infeasible for current advanced tactile sensors. In particular, LightTact enables reliable perception of extremely light contact during interaction with liquids, semi-liquids, and ultra-soft or thin materials, where deformation-based VBTSs and force sensing provide uninformed or weak signals. Moreover, we demonstrate that LightTact’s visual–tactile outputs can be directly interpreted by Vision-Language Models (VLMs), enabling fine-grained reasoning such as resistor value inference for robotic sorting.
@@ II-A Contact Sensing with Vision-Based Tactile Sensors
Vision-based tactile sensors (VBTSs) use internal cameras to image tactile signals with high spatial sensitivity at low cost, but most detect contact indirectly via macroscopic deformation of a compliant surface. Common approaches include tracking painted markers [yamaguchi2016combining, ward2018tactip, alspach2019soft, cui2021hand, zhang2022deltact, zhang2022tac3d], observing magnified reflective patterns [saga2007high, saga2013precise], or measuring luminance variations of reflective layers [yuan2017gelsight, lin20239dtact]. Because they require visible deformation, the surface must be soft, the contacting object relatively harder, and the applied force above a detectable threshold, limiting light contacts with negligible indentation (e.g., water, facial cream, thin films).
To decouple sensing from deformation, several sensors [han2005low, shimonomura2016robotic, zhang2023tirgel, sun2025soft] apply total internal reflection (TIR) within a transparent medium; contact locally frustrates TIR and increases reflected light, highlighting contact regions. However, this strategy is sensitive to ambient illumination: external light, reflections from non-contact portions of the object, and internal LED leakage can enter the camera through non-contact regions, corrupting the intended TIR signal. As a result, these systems often require monochromatic objects or controlled dark environments.
In contrast, LightTact uses an ambient-blocking optical layout that suppresses both external and internal illumination at non-contact regions while capturing only contact-generated scattered light, keeping non-contact pixels near-black in raw images and enabling robust pixel-level contact segmentation across lighting conditions and object appearances.
Similar objectives appear in non-robotic optical devices, such as the Wyman–White fingerprint imager [wyman1965method] and the Drawing Prism [greene1985drawing]. However, these systems target static capture or tracing and assume more relaxed constraints (e.g., controlled lighting and limited object types), often relying on bulky geometries, dark environments, or even a thin liquid layer to remove air gaps—conditions incompatible with dynamic robotic tactile perception in everyday settings. These conditions are incompatible with robotic tactile perception.
@@ III Methodology
Our goal is to design LightTact as a robotic tactile sensor capable of reliable, pixel-level contact segmentation across diverse objects and lighting conditions. To this end, we first introduce a novel optical layout and sensing principle and analyze its theoretical behavior in Section III-A. We then present how LightTact realizes this principle within a compliant, fingertip-scale form factor suitable for robotic integration in Section III-B. Following, we describe the key fabrication steps for constructing LightTact sensors in Section III-C. Finally, we introduce the procedures for sensor calibration and algorithms for contact segmentation in Section III-D.
@@ III-A Optical Layout and Sensing Principle
Figure 2: Optical layout and sensing principle of LightTact. (a) Core components: a transparent medium with a touching and viewing surface, an internal LED, and a camera. (b) External light entering through non-contact regions is rejected by refraction followed by total internal reflection (TIR). (c) Internal LED illumination reflecting from non-contact regions is also rejected by TIR when the LED is placed within the purple area to satisfy $\theta_{it}<\theta_{i}$. (d) At true contact, the air gap is removed and the contacting surface produces diffuse scattering. A subset of these rays reaches the viewing surface with incidence $\theta_{c}$ and refracts toward the camera.
Figure 2 (a) shows the core optical elements of LightTact: a transparent medium that interacts with external objects, an internal LED for illumination, and a camera that observes light exiting the medium. Two surfaces of the medium play distinct optical roles: the touching surface, where contact occurs, and the viewing surface, through which light exits toward the camera. The remaining surfaces are coated matte-black to absorb any light that reaches them. At the touching surface, combining two light sources (external light and internal) with two region types (contact and non-contact) produces three relevant light behaviors for non-luminous objects: (1) external light incident on non-contact regions, (2) internal illumination reflected on non-contact regions, and (3) internal illumination scattered at contact regions.
Our objective is to obtain deformation-independent contact segmentation from the raw image. To achieve this, the optical layout must completely suppress external and internal light at non-contact regions, while transmitting only light generated at true contacts. The key is a side-view imaging layout: the camera observes the medium through a viewing surface that forms an angle $\theta_{tv}$ with the touching surface, rather than being parallel. This non-parallel geometry separates the light paths, so that only contact-generated light can transmit through the viewing surface and reach the camera. Table I summarizes these behaviors, and we detail them below.
External light rejection at non-contact regions
As illustrated in Fig. 2 (b), external rays approach the touching surface from a wide range of directions. The sensing region corresponds to the span between the two depicted ray bundles. When external light enters the medium through non-contact areas, it refracts at angles below the critical angle $\theta_{c}$ due to the air–medium index difference. A portion of these refracted rays then propagate toward the viewing surface. By choosing a sufficiently large wedge angle ($\theta_{tv}>2\theta_{c}$), all such rays arrive at the viewing surface with incidence $\theta_{c}$, causing them to undergo total internal reflection (TIR) and redirecting them toward the black surfaces. Consequently, no external light from non-contact regions reaches the camera.
Internal illumination rejection at non-contact regions
Internal LED illumination should contribute to the image only at true contacts. As shown in Fig. 2 (c), LED rays that reflect specularly off non-contact regions of the touching surface, must also reach the viewing surface with incidence $\theta_{c}$ to ensure TIR. This is guaranteed by constraining the LED position: the LED must lie entirely within the purple admissible region such that all emitted rays satisfy $\theta_{it}<\theta_{i}=\theta_{tv}-\theta_{c}$. Under this placement rule, every specularly reflected LED ray is rejected by TIR at the viewing surface.
@@
Internal Contact Diffuse scattering Refraction $\theta_{tv}<\frac{\pi}{2}+\theta_{c}$
@@
When an object touches the medium, the air gap disappears and the contacting surface produces diffuse scattering of the LED illumination (Fig. 2(d)). This generates a wide angular spread of rays, which can reach the viewing surface with incidence $\theta_{sv}\in\left(\theta_{tv}-\frac{\pi}{2},\frac{\pi}{2}\right)$. To observe the contacting surface, a subset of these scattered rays must refract through the viewing surface and reach the camera. This requires that the incidence range $\left(\theta_{tv}-\frac{\pi}{2},\frac{\pi}{2}\right)$ intersects the transmissible range $\left(-\theta_{c},\theta_{c}\right)$, which is satisfied when $\theta_{tv}<\frac{\pi}{2}+\theta_{c}$. Under this condition, a subset of diffuse-scattering rays transmits within the green regions and enters the camera, revealing the natural local appearance of the object.
Our optical layout enforces $\theta_{tv}>2\theta_{c}$ so that any light entering through non-contact regions arrives at the viewing surface with incidence $\theta_{c}$ and is therefore rejected by TIR, preventing it from reaching the camera. Strictly speaking, $\theta_{tv}>2\theta_{c}$ is not the only feasible configuration: when $\theta_{tv}<2\theta_{c}$, the camera could, in principle, be placed outside the refracted ray field from non-contact regions. In practice, however, this alternative sacrifices compactness and robustness, as it requires a larger camera offset and stricter suppression of stray reflections. We therefore adopt $\theta_{tv}>2\theta_{c}$ as a simple and reliable default; a detailed analysis of the $\theta_{tv}<2\theta_{c}$ case is provided in Appendix VI-A.
@@ III-B Sensor Design
As shown in Fig. 3(a), LightTact is extremely compact, with a fingertip-scale size of only $12\text{mm}\times 18\text{mm}\times 34.5\text{mm}$. Figures 3(b) and 3(c) illustrate the internal assembly, which mirrors the optical system described in Section III-A: a transparent medium (2, 3), an internal LED (5), and a camera (7), all housed within a rigid shell (4), a surrounding black gel layer (1), and lightweight mounts (6, 8). The following paragraphs describe each component in detail.
Figure 3: Design of LightTact. (a) Assembled view. (b) Components of LightTact illustrated in exploded view. (c) Schematic diagram.
Transparent gel and acrylic window. For most robotic applications, the touching surface should be compliant to support safe and stable interaction. Importantly, this compliance is a design preference rather than a sensing requirement, since our optical principle does not rely on macroscopic deformation. However, if the entire transparent medium is soft, the viewing surface would deform under load and introduce image distortion. To avoid this, we construct the medium as a composite: a soft, transparent gel that forms the touching surface, and a rigid acrylic window that defines a deformation-free viewing surface. The two materials have similar refractive indices ($\approx 1.45$), corresponding to a critical angle of $\theta_{c}\approx 43^{\circ}$. To satisfy the geometric conditions in Section III-A, namely $\theta_{tv}>2\theta_{c}\approx 86^{\circ}$ to suppress non-contact light and $\theta_{tv}<90^{\circ}+\theta_{c}\approx 133^{\circ}$ to transmit contact-scattered light, we set the touching and viewing surfaces to be perpendicular.
LED and camera. Internal illumination is provided by a single 2835-SMD LED, mounted at an oblique angle ($\approx 45^{\circ}$) to improve lighting uniformity across the touching surface. Imaging is performed by a compact UVC camera (OV5693) with a $120^{\circ}$ field of view, placed close to the transparent medium to preserve a fingertip-scale form factor. Both components are secured with lightweight mounts that are attached to the sensor shell using M2 screws and heat-set inserts.
Sensor shell. The shell not only holds the optical components at their required poses, but also enforces three key geometric-optical constraints: the transparent gel’s geometry, the LED’s illumination envelope, and the camera’s effective viewing range. These design aspects are detailed below.
@@
Although the shell is printed in black to approximate the ideal light-absorbing boundaries, practical materials still exhibit residual reflectance. As shown in Fig. 3(c), surface A can reflect stray light into the camera through non-contact regions. To mitigate this, we design surface A to be vertical ($\theta_{s}=\frac{\pi}{2}$), reducing the chance that specular reflections couple into the viewing path. For applications that require access to confined spaces, the shell can be made wedge-shaped. We analyze the case’s ($\theta_{s}<\frac{\pi}{2}$) optical consequences in Appendix VI-C.
@@
To prevent the LED from directly illuminating surface A, the shell includes two baffles (surfaces B and C) that bound the emission cone. Increasing baffle height improves shielding but is limited by two factors: (1) if surface B extends above point D, the exposed inclined region becomes directly illuminated and increases non-contact light leakage; and (2) if surface C exceeds point E, it occludes part of the contact-generated scattered light from the left portion of the touching surface.
@@
The shell also apertures the camera so that it observes only the touching surface. Allowing the camera to view outside this region would admit additional stray light (e.g., scattering from the rough top or bottom faces of the acrylic window) and interfere the pixels associated with the touching surface.
Black gel. To maintain compliance across the full contact area, we cast a soft black gel around the perimeter of the transparent gel, creating a smooth transition between the touching surface and the rigid shell. It also covers the top surface of the acrylic window, which is too rigid and rough to serve as part of the touching surface.